A piece of newly purchased industrial equipment costs $916,513.00 and is classified as a seven-year property under MACRS. What is the book value of this piece of equipment at the end of year 3?
Answer #1

Book value of equipment at the end of year 7 is $ 4,00,791.13

Depreciation Schedule:
Year Cost Depreciation rate Depreciation expense Accumulated Depreciation expense Book Value
a b c=a*b d e=a-d
1 $ 9,16,513.00 14.29% $ 1,30,969.71 $ 1,30,969.71 $ 7,85,543.29
2 $ 9,16,513.00 24.49% $ 2,24,454.03 $ 3,55,423.74 $ 5,61,089.26
3 $ 9,16,513.00 17.49% $ 1,60,298.12 $ 5,15,721.87 $ 4,00,791.13